# FLAME: Learning to Navigate with Multimodal LLM in Urban Environments

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated potential in Vision-and-Language Navigation (VLN) tasks, yet current applications face challenges. While LLMs excel in general conversation scenarios, they struggle with specialized navigation tasks, yielding suboptimal performance compared to specialized VLN models. We introduce FLAME (FLAMingo-Architected Embodied Agent), a novel Multimodal LLM-based agent and architecture designed for urban VLN tasks that efficiently handles multiple observations. Our approach implements a three-phase tuning technique for effective adaptation to navigation tasks, including single perception tuning for street view description, multiple perception tuning for trajectory summarization, and end-to-end training on VLN datasets. The augmented datasets are synthesized automatically. Experimental results demonstrate FLAME’s superiority over existing methods, surpassing state-of-the-art methods by a 7.3% increase in task completion rate on Touchdown dataset. This work showcases the potential of Multimodal LLMs (MLLMs) in complex navigation tasks, representing an advancement towards practical applications of MLLMs in embodied AI. Project page: https://flame-sjtu.github.io
@@ 1 Introduction
Figure 1: LLM-based agents excel in conversation but often falter in specialized navigation tasks. Our agent, powered solely by a Multimodal LLM, demonstrates proficiency in navigation skills, adapting to navigation-specific scenarios through targeted finetuning phases.
Large Language Models (LLMs) (Achiam et al. 2023; Touvron et al. 2023) have revolutionized the field of embodied intelligence. Vision-and-Language Navigation (VLN) (Anderson et al. 2018), a fundamental task in embodied AI, challenges agents to navigate to a goal following human instructions in indoor or outdoor environments. This task demands sophisticated abilities in instruction comprehension, environmental understanding, and decision-making, which can be effectively managed by LLMs. Recent approaches have integrated LLMs into VLN methods, either by translating visual data into language (Zhou, Hong, and Wu 2024; Pan et al. 2024; Qiao et al. 2023; Lin et al. 2024a; Chen et al. 2024) or by employing Multimodal LLMs (MLLMs) (Zhang et al. 2024; Zhou et al. 2024) for environmental perception.
However, the incorporation of general-purpose LLMs in VLN still faces critical challenges, primarily stemming from their inherent limitations in navigation-specific scenarios, as Figure 1(a) depicts. For text-only LLMs, translating visual data into language using visual foundation models can lead to information loss (Zhou, Hong, and Wu 2024), resulting in a huge performance gap compared to VLN-specialized models. For Multimodal LLMs, while partially addressing the limitations of text-only LLMs, they often struggle to interact within navigation-specific scenarios. Recent attempts to incorporate MLLMs for navigation focus on integration with down-stream models (Zhou et al. 2024) or intensive training on navigation videos (Zhang et al. 2024), rather than making efficient adaptations to navigation-specific scenario that would harness the MLLMs’ inherent capabilities in handling interleaved textual and visual inputs, leading to suboptimal performance.
@@
To address these challenges, we introduce FLAME (FLAMingo-Architected Embodied Agent), the first MLLM-based agent designed for urban VLN tasks, as shown in Figure 1(b). Based on Flamingo (Alayrac et al. 2022), FLAME operates autoregressively and efficiently handles multiple perceptions without increasing context length, ensuring efficiency in end-to-end training and inference. We propose a three-phase tuning technique to adapt Flamingo model (Alayrac et al. 2022) to navigation tasks using augmented data: 1) Single perception tuning: Learning to describe street views. 2) Multiple perception tuning: Learning to summarize agent trajectories. 3) End-to-End training and evaluation on VLN datasets. To support the first two tuning phases, we utilize GPT-4 (Achiam et al. 2023) to synthesize captions and route summaries for the Touchdown environment (Chen et al. 2019). Additionally, we synthesize navigation rationales for urban VLN datasets (Chen et al. 2019; Schumann and Riezler 2021) to validate FLAME’s reasoning capability (Wei et al. 2022).
Experimental results demonstrate FLAME’s superiority over existing methods on two urban VLN datasets: Touchdown (Chen et al. 2019) and Map2seq (Schumann and Riezler 2021). Our approach significantly outperforms current state-of-the-art (SOTA) methods by 7.3% TC in Touchdown and 3.74% TC in Map2seq, establishing new standards for urban VLN tasks. Our work not only benefits the field of urban VLN but also showcases the potential of MLLMs in navigation within complex environments and tasks.
@@ 2 Related Works
2.1 Vision-and-Language Navigation
Vision-and-Language Navigation (VLN) (Anderson et al. 2018) encompasses indoor (Qi et al. 2020; Krantz et al. 2020) and outdoor scenarios (Chen et al. 2019; Schumann and Riezler 2021), with advancements in multimodal understanding (Tan, Yu, and Bansal 2019; Hong et al. 2021; Zhu et al. 2020) and action strategies (Qi et al. 2020; Chen et al. 2022b), primarily for indoor environments. Traditional VLN agents often lack advanced decision-making skills, prompting the integration of Large Language Models (LLMs) (Lin et al. 2024b, a; Zhan et al. 2024b; Schumann et al. 2023), leveraging the reasoning (Pan et al. 2024; Chen et al. 2024; Lin et al. 2024a) and dialogue capabilities (Qiao et al. 2023; Long et al. 2024) of LLMs. These approaches either convert visual data to text (Zhou, Hong, and Wu 2024) or employ Multimodal LLMs (MLLMs) with intensive training (Zhang et al. 2024). Given the relative underexploration of outdoor VLN, our work addresses this gap by introducing an effectively adapted MLLM-based agent for urban VLN tasks.
2.2 Multimodal Large Language Models
The emergence of Multimodal Large Language Models (MLLMs) (Liu et al. 2024; Wang et al. 2023a; Zhu et al. 2024; Alayrac et al. 2022; Awadalla et al. 2023) has expanded their application to various tasks, including captioning (Li et al. 2023b), grounding (Peng et al. 2023), and general instruction following (Dai et al. 2023; Bai et al. 2023). These models demonstrate multimodal reasoning skills (Lu et al. 2022; Zeng et al. 2023; Lu et al. 2024) in chat conversations, handling single-turn, in-context (Sun et al. 2024), and interleaved text and image inputs (Laurençon et al. 2024a; Jiang et al. 2024). MLLMs have benefited from vision-and-language tuning (Laurençon et al. 2024b), enabling them to process diverse modalities (Zhan et al. 2024a; Wu et al. 2023). However, the general pretraining is insufficient for expert navigation tasks. Our work addresses this gap by adapting an MLLM from general scenarios to navigation through a specialized tuning approach.
2.3 Data Augmentation in Vision-and-Language Navigation
To overcome data scarcity in navigation, various data augmentation techniques (Chen et al. 2022a; Ku et al. 2020; Zhao et al. 2021; Huang et al. 2019b, a) have been proposed. These include utilizing a speaker module (Fried et al. 2018; Dou and Peng 2022) to generate synthetic instructions, leveraging multilingual data (Li, Tan, and Bansal 2022a), incorporating counterfactual information (Wang et al. 2022; Parvaneh et al. 2020; Fu et al. 2020), and altering environments (Li, Tan, and Bansal 2022b; Liu et al. 2021). For outdoor VLN, previous studies have explored training agents on instructions with different styles (Zhu et al. 2021), pretraining on auxiliary tasks (Armitage, Impett, and Sennrich 2023) and using driving videos (Li et al. 2024). However, the use of auxiliary training data to tailor MLLMs for outdoor navigation remains largely unexplored. Our work addresses this gap by using synthetic data to adapt MLLMs for urban VLN tasks.
@@ 3 Method
Figure 2: Overview of FLAME’s navigation process at time step $t_{n}$. The model architecture, based on Flamingo, integrates vision modules for observation processing and decoder blocks for instruction and history handling. The finetuned STRIDED GATED XATTN layers prioritize recent observations in cross-attention computation. At key locations, FLAME can engage in reasoning before decision-making or proceed directly to action selection. The navigation process is autoregressive.
Figure 3: Illustration of the three-phase tuning for navigation and synthetic data generation process. (a) The first phase trains the model on single-perception tasks. The second phase escalates to handling multi-perceptual input. Finally, the model undergoes an end-to-end finetuning. (b) We utilize LLMs to generate street view captions and route summaries to aid the training of the first two phase. (c) We further synthesize rationales for to validate the reasoning capability of FLAME.
@@
3.1 FLAME Architecture
FLAME builds upon the Flamingo architecture (Alayrac et al. 2022), leveraging cross-attention for processing of visual and textual inputs, without extending context length and improves efficiency. We introduce two key adaptations to Flamingo for urban VLN:
@@ 3.1 FLAME Architecture
where CMA denotes cross-modal attention, $W_{Q}$, $W_{K}^{v}$, $W_{V}^{v}$ are learnable parameters, and $x_{j}^{v}$ refers to the $j$-th visual token. This approach is aimed at prioritizing recent observations, thereby augmenting the system’s proficiency in identifying and responding to significant features in the environment dynamically.
@@
As Figure 2(c) shows, FLAME operates autoregressively, predicting actions based on the initial instruction $I$, current observation $O_{t}$, history of observations $O_{\leq t-1}$ and previous actions. By default, at each timestep $t$, the next action $a_{t}$ is obtained by:
@@
Optionally, to make agent’s thought process transparent and understandable, FLAME generate rationales at key locations (e.g., intersections) before action prediction, similar to ReAct (Yao et al. 2023). Let $R_{t}$ denote the generated rationale, the next action can be obtained by:
@@
3.2 Three-Phase Tuning for Navigation
To adapt Flamingo for urban VLN tasks, we propose a three-phase tuning paradigm, as depicted in Figure 3(a). This approach progressively builds the model’s capabilities from understanding environment to complex decision-making.
@@ 3.2 Three-Phase Tuning for Navigation
In the first phase, we train FLAME on a street view captioning task to develop its feature recognition abilities. Given a dataset $\mathcal{D}_{p1}=\{\tau^{(i)}\}_{i=1}^{N}$, where each instance $\tau^{(i)}=\{(P^{(i)},O^{(i)},c^{(i)})\}$ consists of a task prompt $P^{(i)}$, an observation $O^{(i)}$, and a ground truth caption $c^{(i)}$, the training objective is defined as:
@@
The second phase focuses on synthesizing information from sequential observations, crucial for navigation understanding. Using a trajectory summary dataset $\mathcal{D}_{p2}=\{\tau^{(i)}\}_{i=1}^{N}$, where $\tau^{(i)}=\{(P^{(i)},O^{(i)}_{1},...,O^{(i)}_{T^{(i)}},s^{(i)})\}$, we minimize:
$\mathcal{L}_{p2}=-\sum_{i=1}^{N}\log p(s^{(i)}|P^{(i)},O^{(i)}_{1},...,O^{(i)}_{T^{(i)}};\theta).$ (5)
Here, $s^{(i)}$ is the ground truth summary, and $T^{(i)}$ is the number of images in the $i$-th instance.
@@
The final phase finetunes FLAME on the VLN dataset $\mathcal{D}_{vln}=\{\tau^{(i)}\}_{i=1}^{N}$, where each instance $\tau^{(i)}=\{(I^{(i)},O^{(i)}_{1},a^{(i)}_{1},...,O^{(i)}_{T^{(i)}},a^{(i)}_{T^{(i)}})\}$ contains an instruction $I^{(i)}$, observations $O^{(i)}_{t}$, and ground truth actions $a^{(i)}_{t}$. We minimize:
$\mathcal{L}_{vln}=-\sum_{i=1}^{N}\sum_{t=1}^{T^{(i)}}\log p(a^{(i)}_{t}|I^{(i)},O^{(i)}_{\leq t},a^{(i)}_{\leq t-1};\theta).$ (6)
This multi-phase approach progressively builds FLAME’s capabilities, from basic urban environment understanding to complex decision-making, enabling it to effectively process navigational instructions and environmental cues.
3.3 Synthetic Data Generation
@@
We focus on generating captions for street views at key locations, which are crucial for successful navigation. These images are processed by GPT-4V (Achiam et al. 2023) to produce corresponding descriptions. To ensure diversity, we randomly select meta-prompts from a curated list, guiding the caption generation process.
Trajectory Summary Generation
We construct a comprehensive knowledge graph of landmarks by combining places of interests from OpenStreetMap with additional map information, following the methodology of (Schumann and Riezler 2021). Using this graph as input, we employ GPT-4 to generate detailed route summaries between key locations.
@@
To validate reasoning capabilities of our model, we generate synthetic rationales for the VLN datasets. The process involves segmenting each VLN trajectory into sub-routes between key locations, then retrieving the corresponding image caption and sub-route summary for each key location. We then utilize GPT-4 to generate a synthetic rationale at each key location, based on the retrieved information. To maintain data quality, we discard instances containing rationales that contradict ground truth actions. This results in a altered subset of the VLN dataset augmented with synthetic rationales, which facilitates end-to-end training with reasoning capabilities.
@@
Table 1: Comparison with state-of-the-art models on Touchdown and Map2seq datasets. Models denoted by (*) utilize image features consistent with FLAME’s implementation. Bold values indicate best performance.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated potential in Vision-and-Language Navigation (VLN) tasks, yet current applications face challenges. While LLMs excel in general conversation scenarios, they struggle with specialized navigation tasks, yielding suboptimal performance compared to specialized VLN models. We introduce FLAME (FLAMingo-Architected Embodied Agent), a novel Multimodal LLM-based agent and architecture designed for urban VLN tasks that efficiently handles multiple observations. Our approach implements a three-phase tuning technique for effective adaptation to navigation tasks, including single perception tuning for street view description, multiple perception tuning for route summarization, and end-to-end training on VLN datasets. The augmented datasets are synthesized automatically. Experimental results demonstrate FLAME’s superiority over existing methods, surpassing state-of-the-art methods by a 7.3% increase in task completion on Touchdown dataset. This work showcases the potential of Multimodal LLMs (MLLMs) in complex navigation tasks, representing an advancement towards applications of MLLMs in the field of embodied intelligence.
Codehttps://github.com/xyz9911/FLAME
@@ 1 Introduction
Figure 1: LLM-based agents excel in conversation but often falter in specialized navigation tasks. Our agent, powered solely by a Multimodal LLM, demonstrates proficiency in navigation skills, efficiently adapting to navigation-specific scenarios through targeted finetuning phases.
Large Language Models (LLMs) (Achiam et al. 2023; Touvron et al. 2023) have revolutionized the field of embodied intelligence. Vision-and-Language Navigation (VLN) (Anderson et al. 2018), a fundamental task in embodied AI, challenges agents to navigate to a goal following human instructions in indoor or outdoor environments. This task demands sophisticated abilities in instruction comprehension, environmental understanding, and decision-making, which can be effectively managed by LLMs. Recent approaches have integrated LLMs into VLN methods, either by translating visual data into language (Zhou, Hong, and Wu 2024; Qiao et al. 2023; Chen et al. 2024) or by employing Multimodal LLMs (MLLMs) (Zhang et al. 2024; Zhou et al. 2024) for environmental perception.
However, the incorporation of general-purpose LLMs in VLN still faces critical challenges, primarily stemming from their inherent limitations in navigation-specific scenarios, as Figure 1(a) depicts. For text-only LLMs, translating visual data into language using visual foundation models can lead to information loss (Zhou, Hong, and Wu 2024), resulting in a performance gap compared to VLN-specialized models. For Multimodal LLMs, while partially addressing the limitations of text-only LLMs, they often struggle to interact within navigation-specific scenarios. Recent attempts to incorporate MLLMs for navigation may have constrained their capabilities by either using them as auxiliary components to traditional VLN models (Zhou et al. 2024) or by processing observations through video tokens (Zhang et al. 2024), which require multiple forward passes per trajectory and significant computational resources in training. These approaches may not efficiently adapt MLLMs to navigation-specific scenarios, which limits their ability to leverage inherent capabilities for handling interleaved textual and visual inputs and affects overall performance.
@@
To address these challenges, we introduce FLAME (FLAMingo-Architected Embodied Agent), the first MLLM-based agent designed for urban VLN tasks, as shown in Figure 1(b). Based on Flamingo (Alayrac et al. 2022), FLAME operates autoregressively and efficiently handles multiple perceptions without increasing context length, ensuring efficiency in end-to-end training and inference. We propose a three-phase tuning technique to adapt Flamingo model (Alayrac et al. 2022) to navigation tasks using augmented data: 1) Single perception tuning: Learning to describe street views. 2) Multiple perception tuning: Learning to summarize agent navigation. 3) End-to-End training and evaluation on VLN datasets. To support the first two tuning phases, we utilize GPT-4 (Achiam et al. 2023) to synthesize captions and route summaries for the Touchdown environment (Chen et al. 2019). Additionally, we synthesize navigation rationales for urban VLN datasets (Chen et al. 2019; Schumann and Riezler 2021) to validate FLAME’s reasoning capability (Wei et al. 2022).
Our agent achieves remarkable computational efficiency, completing the entire training process in just 14 hours on a single A100 GPU. Experimental results demonstrate FLAME’s superiority over existing methods on two urban VLN datasets: Touchdown (Chen et al. 2019) and Map2seq (Schumann and Riezler 2021). Our approach significantly outperforms current state-of-the-art (SOTA) methods by 7.3% Task Completion (TC) in Touchdown and 3.74% TC in Map2seq. Our work not only benefits the field of VLN but also showcases the potential of MLLMs in navigation within complex environments and tasks.
@@ 2 Related Works
Vision-and-Language Navigation
Vision-and-Language Navigation (VLN) (Anderson et al. 2018) encompasses indoor (Qi et al. 2020; Ku et al. 2020) and outdoor scenarios (Chen et al. 2019; Schumann and Riezler 2021), with most advancements focusing on indoor environments. Traditional VLN agents often lack advanced decision-making skills, prompting the integration of Large Language Models (LLMs) (Lin et al. 2024; Schumann et al. 2024), leveraging the reasoning (Chen et al. 2024) and dialogue capabilities (Qiao et al. 2023; Long et al. 2024) of LLMs. These approaches either convert visual data to text (Zhou, Hong, and Wu 2024) or employ Multimodal LLMs (MLLMs) with intensive training (Zhang et al. 2024; Zheng et al. 2024). Given the relative underexploration of outdoor VLN, our work addresses this gap by introducing an effectively adapted MLLM-based agent for urban VLN tasks.
Multimodal Large Language Models
The emergence of Multimodal Large Language Models (MLLMs) (Liu et al. 2024; Alayrac et al. 2022; Awadalla et al. 2023) has expanded their application to various tasks, including captioning (Li et al. 2023b) and general instruction following (Dai et al. 2023). These models demonstrate multimodal reasoning skills (Lu et al. 2022) in chat conversations, handling single-turn, in-context (Sun et al. 2024), and interleaved text and image inputs (Laurençon et al. 2024). MLLMs have benefited from vision-and-language tuning, enabling them to process diverse modalities (Wu et al. 2024). However, the general pretraining is insufficient for expert navigation tasks. Our work addresses this gap by adapting an MLLM from general scenarios to navigation through a specialized tuning approach.
Data Augmentation in Vision-and-Language Navigation
To overcome data scarcity in navigation, various data augmentation techniques (Zhao et al. 2021; Huang et al. 2019) have been proposed. These include utilizing a speaker module (Fried et al. 2018; Dou and Peng 2022) to generate synthetic instructions, leveraging multilingual data (Li, Tan, and Bansal 2022a), incorporating counterfactual information (Parvaneh et al. 2020; Fu et al. 2020), and altering environments (Li, Tan, and Bansal 2022b; Liu et al. 2021). For outdoor VLN, previous studies have explored training agents on instructions with different styles (Zhu et al. 2021), pretraining on auxiliary tasks (Armitage, Impett, and Sennrich 2023) and using driving videos (Li et al. 2024). However, the use of auxiliary training data to tailor MLLMs for outdoor navigation remains largely unexplored. Our work addresses this gap by using synthetic data to adapt MLLMs for urban VLN tasks.
@@ 3 Method
Figure 2: Overview of FLAME’s navigation process at time step $t_{n}$. The architecture, based on Flamingo, integrates vision modules for observation processing and decoder blocks for instruction and history handling. The finetuned STRIDED GATED XATTN layers prioritize recent observations in cross-attention computation. At key locations (intersections), FLAME can engage in reasoning before decision-making or proceed directly to action selection. The navigation process is autoregressive.
Figure 3: Illustration of the three-phase tuning for navigation and synthetic data generation process. (a) The first phase trains the model on single-perception tasks. The second phase escalates to handling multi-perceptual input. Finally, the model undergoes an end-to-end finetuning. (b) We utilize LLMs to generate street view captions, route summaries and simple instructions to aid the training of the first two phase. (c) We further synthesize rationales to validate the reasoning capability of FLAME.
@@
FLAME Architecture
FLAME builds upon the Flamingo architecture (Alayrac et al. 2022), leveraging cross-attention to process visual and textual inputs without extending context length, while ensuring single-pass fine-tuning per trajectory. We introduce two key adaptations to Flamingo for urban VLN:
@@
where CMA denotes cross-modal attention, $W_{Q}$, $W_{K}^{v}$, $W_{V}^{v}$ are learnable parameters, and $x_{j}^{v}$ refers to the $j$-th visual token. This approach is aimed at prioritizing recent observations, thereby augmenting the system’s proficiency in identifying significant features in the dynamic environment.
@@
As Figure 2(c) shows, FLAME predicts actions based on the instruction $I$, current observation $O_{t}$, history of observations $O_{\leq t-1}$ and previous actions. By default, the next action $a_{t}$ is obtained by:
@@
Optionally, to make agent’s thought process transparent and understandable, FLAME generate rationales at key locations (i.e., intersections) before action prediction, similar to ReAct (Yao et al. 2023). Let $R_{t}$ denote the generated rationale, the next action can be obtained by:
@@
Three-Phase Tuning for Navigation
To adapt Flamingo for urban VLN tasks, we propose a three-phase tuning paradigm, as depicted in Figure 3(a).
@@
In the first phase, we train the model on a street view captioning task to strengthen its feature recognition abilities. Given a dataset $\mathcal{D}_{p1}=\{\tau^{(i)}\}_{i=1}^{N}$, where $\tau^{(i)}=\{(P^{(i)},O^{(i)},c^{(i)})\}$ consists of a caption prompt $P^{(i)}$, an observation $O^{(i)}$, and a ground truth caption $c^{(i)}$, the training objective is:
@@
The second phase focuses on synthesizing sequential observations and performing actions. Using a dataset $\mathcal{D}_{p2}=\{\tau^{(i)}\}_{i=1}^{N}$, where each instance $\tau^{(i)}=\{(I^{(i)},O^{(i)}_{1},a^{(i)}_{1},...,O^{(i)}_{T^{(i)}},a^{(i)}_{T^{(i)}},s^{(i)})\}$ consists of a simple instruction $I^{(i)}$, observations $O^{(i)}_{\leq T^{(i)}}$, actions $a^{(i)}_{\leq T^{(i)}}$ and a ground truth route summary $s^{(i)}$. We supervise two objectives:
$\mathcal{L}_{sum}^{(i)}=\log p(s^{(i)}|I^{(i)},O_{\leq T^{(i)}}^{(i)};\theta),$ (5)
$\mathcal{L}_{act}^{(i)}=\sum_{t=1}^{T^{(i)}}\log p(a_{t}^{(i)}|I^{(i)},O_{\leq t}^{(i)},a_{\leq t-1}^{(i)};\theta),$ (6)
where $\mathcal{L}_{sum}^{(i)}$ is the route summarization loss and $\mathcal{L}_{act}^{(i)}$ is the imitation loss. The total loss for the second phase is:
$\mathcal{L}_{p2}=-\sum_{i=1}^{N}\Big{(}\mathcal{L}_{act}^{(i)}+\mathcal{L}_{sum}^{(i)}\Big{)}.$ (7)
@@
Finally, FLAME is finetuned on a VLN dataset $\mathcal{D}_{nav}=\{\tau^{(i)}\}_{i=1}^{N}$, where $\tau^{(i)}=\{(I^{(i)},O^{(i)}_{1},a^{(i)}_{1},...,O^{(i)}_{T^{(i)}},a^{(i)}_{T^{(i)}})\}$. We minimize:
$\mathcal{L}_{nav}=-\sum_{i=1}^{N}\sum_{t=1}^{T^{(i)}}\log p(a^{(i)}_{t}|I^{(i)},O^{(i)}_{\leq t},a^{(i)}_{\leq t-1};\theta).$ (8)
This multi-phased approach progressively builds the model’s capabilities, from environment understanding to complex decision-making.
Synthetic Data Generation
@@
We focus on generating captions for street views at key locations, which are crucial for navigation. These images are processed by GPT-4V (Achiam et al. 2023) to produce descriptions. To ensure diversity, we randomly select meta-prompts from a handcrafted prompt list, guiding the caption generation process. The prompt list contains specialized prompts refined manually to elicit diverse street view descriptions from GPT-4V.
Route Summary Generation
We construct a comprehensive knowledge graph of landmarks by combining places of interests from OpenStreetMap with additional map information, following the methodology of (Schumann and Riezler 2021). Using this graph as input, we employ GPT-4 to generate detailed route summaries and simple navigation instructions between key locations.
@@
To validate reasoning capabilities of our agent, we generate synthetic rationales for the VLN datasets. The process involves segmenting each VLN trajectory into sub-routes, then retrieving the corresponding image caption and summary for each route. We then utilize GPT-4 to generate a synthetic rationale at each key location, based on the retrieved information. To maintain data quality, we discard instances with flawed rationales. This results in a altered subset of the VLN dataset augmented with synthetic rationales, which facilitates end-to-end training with reasoning capabilities.
@@
Table 1: Comparison with state-of-the-art models on Touchdown and Map2seq datasets. Models denoted by (*) utilize visual context consistent with FLAME’s implementation. Bold values indicate best performance.